S is the sum of the first 2n terms of a GP, if the sum of the alternate terms of the same gp is S/4, find the common ratio of the GP
Answer by stanbon(75887)

S is the sum of the first 2n terms of a GP, if the sum of the alternate terms of the same gp is S/4, find the common ratio of the GP
----------------------
S(2n) = a[(r^(2n)-1)/(r-1)] = S
S(n) = a[(r^n-1)/(r-1)] = S/4
================
Divide S(2n)/S(n) to get:
(r^(2n)-1)/(r^n-1) = 4
------
r^(2n)-1 = 4r^n - 4
------
(r^n)^2-1-4r^n+4 = 0
(r^n)^2 - 4r^n + 3 = 0
(r^n-1)(r^n-3) = 0
r^n = 1 or r^n = 3
=====================
Cheers,
Stan H.
====================